OpenStudy (anonymous):

Twenty stamps cost $8.40. How much does it cost to buy 5 stamps?

OpenStudy (anonymous):

divide the 8.40 by 20. which is 0.42. then multiply by 5 which is $2.10

OpenStudy (anonymous):

20 / 4 = 5 so divide 8.40 by 4 and you'll get your answer

OpenStudy (anonymous):

or you can take the long route Link proposed you with.

OpenStudy (anonymous):

thxs

OpenStudy (texaschic101):

or set up a proportion.. 20/8.40 = 5/x (20 stamps to 8.40 = 5 stamps to x) cross multiply 20x = (8.40)(5) just finish the rest and solve for x and that will be your answer